# Bandit and Delayed Feedback in Online Structured Prediction

### v1

@@ 3.2 Online Gradient Descent
The algorithm in this section uses the adaptive Online Gradient Descent (OGD, Streeter & McMahan 2010) as $\mathsf{ALG}$. OGD updates $\bm{W}_{t}$ to $\bm{W}_{t+1}$ using a gradient $\bm{G}_{t}$ and learning rate $\eta_{t}$ by $\bm{W}_{t+1}\leftarrow\Pi_{\mathcal{W}}\left(\bm{W}_{t}-\eta_{t}\bm{G}_{t}\right),$ where $\Pi_{\mathcal{W}}(\bm{Z})=\operatorname*{arg\,min}_{\bm{X}\in\mathcal{W}}\|\bm{X}-\bm{Z}\|_{\mathrm{F}}$. OGD with appropriately chosen learning rate $\eta_{t}$ achieves the following bound:
Lemma 3.3 (Orabona 2019, Theorem 4.14).
Suppose that we set the learning rate to $\eta_{t}=\frac{B}{\sqrt{2\sum_{i=1}^{t}\|\bm{G}_{i}\|_{\mathrm{F}}^{2}}}$ and do not update on rounds when $\bm{G}_{t}$ is the all-zero matrix. Then, for any $\bm{U}\in\mathcal{W}$, OGD achieves $\sum_{t=1}^{T}{\langle\bm{G}_{t},\bm{W}_{t}-\bm{U}\rangle}\leq\sqrt{2}B\sqrt{\sum_{t=1}^{T}{\|\bm{G}_{t}\|_{\mathrm{F}}^{2}}}.$
3.3 $O(\sqrt{KT})$ Regret Algorithm
Here, we present an algorithm that achieves a regret upper bound of $O(\sqrt{KT})$.
Algorithm based on inverse-weighted estimator
In the bandit setting, the true output $\bm{y}_{t}$ is not observed, and thus it is necessary to estimate the gradient required for updating $\bm{W}_{t}$. To deal with this, we use the following inverse-weighted gradient estimator:
$\widetilde{\bm{G}}_{t}\coloneqq\frac{\mathds{1}[\widehat{\bm{y}}_{t}=\bm{y}_{t}]}{p_{t}(\bm{y}_{t})}\bm{G}_{t},$ (4)
where we recall that $\bm{G}_{t}=\nabla S_{t}(\bm{W}_{t})=(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t})\bm{x}_{t}^{\top}$. Note that $\widetilde{\bm{G}}_{t}$ is unbiased, i.e., $\mathbb{E}\big{[}\widetilde{\bm{G}}_{t}\big{]}=\bm{G}_{t}$. We use RDUE with $q=B\sqrt{K/T}$ as the decoding function (assuming $T\geq B^{2}K$ for simplicity). For $\mathsf{ALG}$, we employ the adaptive OGD in Section 3.2 with the learning rate of $\eta_{t}=\frac{B}{\sqrt{2\sum_{i=1}^{t}\|\widetilde{\bm{G}}_{i}\|_{\mathrm{F}}^{2}}}.$
@@ 3.3 $O(\sqrt{KT})$ Regret Algorithm
The above algorithm achieves the following regret bound:
@@
The regret of the above algorithm is bounded by $\mathbb{E}[\mathcal{R}_{T}]\leq\left(\frac{b}{2a}+1\right)B\sqrt{KT}.$
The $O(\sqrt{KT})$ bound has an optimal dependency on $T$ and matches the $\sqrt{T}$ lower bound in the special case of online multiclass classification with bandit feedback (Van der Hoeven et al., 2021, Corollary 1). Furthermore, our bound improves the existing $O(\sqrt{KT})$ bound by Van der Hoeven (2020) by a factor of $\sqrt{K}$. Note that, due to differences in the target loss function, our result is not directly comparable to the $\sqrt{KT}$ bound in Van der Hoeven et al. (2021). A more detailed discussion can be found in Appendix C.
@@
From the convexity of $S_{t}$ and the unbiasedness of $\widetilde{\bm{G}}_{t}$, we have $\mathbb{E}\left[\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}\right]\leq\mathbb{E}\left[\sum_{t=1}^{T}{\langle\bm{G}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]=\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widetilde{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right].$ From Lemma 3.3, this is further upper bounded as $\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widetilde{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]\leq\sqrt{2}B\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}}\right]}\leq B\sqrt{\frac{2bK}{q}\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]},$ where in the first inequality we used Jensen’s inequality and in the last inequality we used $\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]=\frac{\|\bm{G}_{t}\|_{\mathrm{F}}^{2}}{p_{t}(\bm{y}_{t})}\leq\frac{K}{q}\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq\frac{bK}{q}S_{t}(\bm{W}_{t}),$ which follows from $p_{t}(\bm{y})\geq K/q$ and (2). Therefore, from Assumption 3.2, we have $\mathbb{E}[\mathcal{R}_{T}]\leq\mathbb{E}\left[\sum_{t=1}^{T}{\left((1-a)S_{t}(\bm{W}_{t})-S_{t}(\bm{U})\right)}\right]+qT\leq B\sqrt{\frac{2bK}{q}\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]}-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT\leq\frac{bB^{2}K}{2aq}+qT,$ where the last inequality follows from $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$. Finally, substituting $q=B\sqrt{K/T}$ into the last inequality yields the desired bound. ∎
We can also prove the following high-probability bound:
@@
Let $\delta\in(0,1)$. Then with probability at least $1-\delta$, the same algorithm as in Theorem 3.5, but with a different choice of $q$, achieves the regret bound of $\mathcal{R}_{T}=O\Big{(}\sqrt{KT\log(1/\delta)}+\log(1/\delta)\Big{)}$, where we omit the dependencies on parameters other than $K$, $T$, and $\delta$.
A more precise statement and proof of this theorem are provided in Section D.2. To prove this theorem, we follow the analysis of Theorem 3.5 and use Bernstein’s inequality. To address the challenges posed by the randomness introduced by bandit feedback, we adopt an approach similar to that used by Van der Hoeven et al. (2021), and arguably, we have successfully simplified their analysis.
3.4 $O(T^{2/3})$ Regret Algorithm
While the $O(\sqrt{KT})$ regret bound given above is desirable in terms of the dependence on $T$, the dependence on $K=\lvert\mathcal{Y}\rvert$ is undesirable for general structured prediction. In fact, we have $K=\binom{d}{m}$ in multilabel classification with $m$ correct labels and $K=m!$ in ranking with $m$ items. To address this issue, we present an algorithm that significantly improves the dependence on $K$ when the target loss function belongs to a special class of SELF satisfying the following assumptions:
@@ 3.4 $O(T^{2/3})$ Regret Algorithm
(i) $\bm{V}$ is invertible, and $\bm{b}$ and $c(\cdot)$ are known and non-negative. (ii) Let $\bm{Q}=\mathbb{E}_{\bm{y}\sim\mu}[\bm{y}\bm{y}^{\top}]$, where $\mu$ is the uniform distribution over $\mathcal{Y}$. At least one of the following two conditions holds: (ii-a) $\bm{Q}$ is invertible, or (ii-b) for any $\bm{y}\in\mathcal{Y}$, $\bm{V}\bm{y}$ lies in the linear subspace spanned by vectors in $\mathcal{Y}$. (iii) For some $\omega>0$, it holds that
$\operatorname*{tr}\left(\bm{V}^{-1}\bm{Q}^{+}(\bm{V}^{-1})^{\top}\right)\leq\omega.$
The first condition is true in the examples in Section 2.4, assuming that the number of correct labels is fixed in multilabel classification. The second one is satisfied if $\mathcal{Y}$ consists of $d$ linearly independent vectors or $\bm{V}$ is proportional to the identity matrix; either is true in those examples. It is also not difficult to derive bounds on $\omega$ in those examples.
Algorithm based on pseudo-inverse matrix estimator
As in the case of Section 3.3, we begin by providing a method to estimate the gradient. Define $\bm{P}_{t}\coloneqq\mathbb{E}_{\bm{y}\sim p_{t}}[\bm{y}\bm{y}^{\top}]$. Then, we define the estimator $\widetilde{\bm{y}}_{t}$ of $\bm{y}_{t}$ by
$\widetilde{\bm{y}}_{t}\coloneqq{\bm{V}}^{-1}\bm{P}_{t}^{+}\widehat{\bm{y}}_{t}\langle\widehat{\bm{y}}_{t},\bm{V}\bm{y}_{t}\rangle,$
where $\bm{P}_{t}^{+}$ is the Moore–Penrose pseudo-inverse matrix of $\bm{P}_{t}$. It is important to note that, given that $\bm{b}$ and $c(\cdot)$ are known, $\langle\widehat{\bm{y}}_{t},\bm{V}\bm{y}_{t}\rangle=L_{t}(\widehat{\bm{y}}_{t})-\langle\widehat{\bm{y}}_{t},\bm{b}\rangle-c(\bm{y}_{t})$ can be computed. Note that $\widetilde{\bm{y}}_{t}$ is unbiased, i.e., $\mathbb{E}_{t}\left[{\widetilde{\bm{y}}_{t}}\right]=\bm{y}_{t}$ from the first requirement of Assumption 3.7.
Using this $\widetilde{\bm{y}}_{t}$, we define the gradient estimator $\widetilde{\bm{G}}_{t}$ by
$\widetilde{\bm{G}}_{t}\coloneqq\left(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\widetilde{\bm{y}}_{t}\right)\bm{x}_{t}^{\top},$ (5)
whose expectation is $\mathbb{E}\big{[}\widetilde{\bm{G}}_{t}\big{]}=\bm{G}_{t}.$ Our estimator is based on the estimators used in adversarial linear bandits and adversarial combinatorial full-bandits (Dani et al., 2007; Abernethy et al., 2008; Cesa-Bianchi & Lugosi, 2012).
We use RDUE with $q=\left(\frac{4\omega B^{2}C_{x}^{2}}{T}\right)^{1/3}$ as the decoding function (assuming $T\geq 4\omega B^{2}C_{x}^{2}$ for simplicity). For updating $\bm{W}_{t}$, we employ the adaptive OGD in Section 3.2 as $\mathsf{ALG}$ with the learning rate of $\eta_{t}=\frac{B}{\sqrt{2\sum_{i=1}^{t}\|\widetilde{\bm{G}}_{i}\|_{\mathrm{F}}^{2}}}.$
Regret bounds
The above algorithm achieves the following regret bound, which does not directly depend on $K$:
Theorem 3.8.
The above algorithm achieves $\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\big{(}\omega^{1/3}\left(BC_{x}T\right)^{2/3}\big{)}.$
The proof can be found in Section D.3. By using the estimator based on the pseudo-inverse matrix, we can upper bound the second moment of the gradient estimator $\widetilde{\bm{G}}_{t}$ without $K$, which allows us to establish the improved regret bound that does not explicitly depend on $K$.
The regret bound in Theorem 3.8 yields the different bounds on each problem setup as follows:
Corollary 3.9.
The above algorithm achieves $\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\left((BC_{x}dT)^{2/3}\right)$ in multiclass classification with the 0-1 loss, $\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\left((d^{5}/m(d-m))^{1/3}(BC_{x}T)^{2/3}\right)$ in multilabel classification with $m$ correct labels and the Hamming loss, and $\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\left(m^{5/3}(BC_{x}T)^{2/3}\right)$ in ranking with the number of items $m$ and the Hamming loss.
The proof of Corollary 3.9 is deferred to Section D.4. The bound for multilabel classification with $m$ correct labels significantly improved the $O(\sqrt{KT})$ bound in Section 3.3, which has a dependency of $\sqrt{\binom{d}{m}}$, and the bound for ranking significantly improved the $O(\sqrt{KT})$ bound in Section 3.3, which has a dependency of $\sqrt{m!}$.
4 Delayed Full-Information Feedback
This section discusses online structured prediction with delayed full-information feedback and provides an algorithm that achieves a surrogate regret bound of $O(D^{2/3}T^{1/3})$, a better bound than $O(\sqrt{DT})$ that can be achieved with a standard OCO algorithm under delayed feedback (Joulani et al., 2013). Below, we make the following assumption.
Assumption 4.1.
There exists a constant $a\in(0,1)$ which satisfies
$\mathbb{E}_{t}\left[{L_{t}(\widehat{\bm{y}}_{t})}\right]\leq(1-a)S_{t}(\bm{W}_{t}).$
@@ 4 Delayed Full-Information Feedback
For updating $\bm{W}_{t}$, we employ the Optimistic Delayed Adaptive FTRL (ODAFTRL) algorithm proposed by Flaspohler et al. (2021). In ODAFTRL, given a gradient $\bm{G}_{t}$ at round $t$, $\bm{W}_{t}$ is updated as
$\bm{W}_{t+1}=\operatorname*{arg\,min}_{\bm{W}\in\mathcal{W}}\left\{\sum_{i=1}^{t-D}\left\langle\bm{G}_{i},\bm{W}\right\rangle+\frac{\lambda_{t}\|\bm{W}\|_{\mathrm{F}}^{2}}{2}\right\},$
where $\lambda_{t}\geq 0$ is the regularization parameter. Due to space constraints, the details of the algorithm are provided in Section E.1. By updating $\lambda_{t}$ using an AdaHedge-type algorithm called AdaHedgeD, ODAFTRL achieves the following regret upper bound:
Lemma 4.2 (Informal version of Flaspohler et al. 2021, Theorem 12).
Consider the setting with delayed full-information feedback. Then, for any $\bm{U}\in\mathcal{W}$, ODAFTRL with the AdaHedgeD update of $\lambda_{t}$ achieves $\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})\!-\!S_{t}(\bm{U}))}\leq\sum_{t=1}^{T}{\langle\bm{G}_{t},\bm{W}_{t}-\bm{U}\rangle}=O\Big{(}\sqrt{\sum_{t=1}^{T}{(\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\!+\!D\|\bm{G}_{t}\|_{\mathrm{F}})}}\Big{)}.$
@@
The algorithm described above achieves the following bound:
@@
The above algorithm achieves $\mathbb{E}[\mathcal{R}_{T}]=O(D^{2/3}T^{1/3}).$
Here, we provide a proof sketch; the complete proof can be found in Section E.2.

### v2

@@ 3.2 Online gradient descent
We use the adaptive Online Gradient Descent (OGD) [36] as $\mathsf{ALG}$, which we apply to surrogate loss $S_{t}$. OGD updates $\bm{W}_{t}$ to $\bm{W}_{t+1}$ by using the gradient $\bm{G}_{t}=\nabla S_{t}(\bm{W}_{t})$ and learning rate $\eta_{t}$ as $\bm{W}_{t+1}\leftarrow\Pi_{\mathcal{W}}\left(\bm{W}_{t}-\eta_{t}\bm{G}_{t}\right),$ where $\Pi_{\mathcal{W}}(\bm{Z})=\operatorname*{arg\,min}_{\bm{X}\in\mathcal{W}}\|\bm{X}-\bm{Z}\|_{\mathrm{F}}$. OGD achieves the following bound:
Lemma 3.3 (e.g., [31, Theorem 4.14]).
Let $\eta_{t}={B}/{\sqrt{2\sum_{i=1}^{t}\|\bm{G}_{i}\|_{\mathrm{F}}^{2}}}$. Then, adaptive OGD achieves $\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}\leq\sum_{t=1}^{T}{\langle\bm{G}_{t},\bm{W}_{t}-\bm{U}\rangle}\leq\sqrt{2}B\sqrt{\sum_{t=1}^{T}{\|\bm{G}_{t}\|_{\mathrm{F}}^{2}}}$ for any $\bm{U}\in\mathcal{W}$.
3.3 Algorithm based on inverse-weighted gradient estimator with $O(\sqrt{KT})$ regret
We present an algorithm that achieves a surrogate regret upper bound of $O(\sqrt{KT})$.
Algorithm based on inverse-weighted gradient estimator
In the bandit setting, the true output $\bm{y}_{t}$ is not observed, and thus we need to estimate the gradient of $S_{t}(\bm{W}_{t})$ required for updating $\bm{W}_{t}$. To do this, we use the inverse-weighted gradient estimator $\widehat{\bm{G}}_{t}\coloneqq\frac{\mathds{1}[\widehat{\bm{y}}_{t}=\bm{y}_{t}]}{p_{t}(\bm{y}_{t})}\bm{G}_{t},$ where $\bm{G}_{t}=\nabla S_{t}(\bm{W}_{t})=(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t})\bm{x}_{t}^{\top}$. Note that $\widehat{\bm{G}}_{t}$ is unbiased, i.e., $\mathbb{E}\big{[}\widehat{\bm{G}}_{t}\big{]}=\bm{G}_{t}$. We use RDUE with $q=B\sqrt{K/T}$ as the decoding function (assuming $T\geq B^{2}K$ for simplicity). For $\mathsf{ALG}$, we employ the adaptive OGD in Section 3.2 with the learning rate of $\eta_{t}={B}/{\sqrt{2\sum_{i=1}^{t}\|\widehat{\bm{G}}_{i}\|_{\mathrm{F}}^{2}}}.$
@@ 3.3 Algorithm based on inverse-weighted gradient estimator with $O(\sqrt{KT})$ regret
The above algorithm achieves the following surrogate regret bound:
@@
The above algorithm achieves the surrogate regret of $\mathbb{E}[\mathcal{R}_{T}]\leq\left(\frac{b}{2a}+1\right)B\sqrt{KT}.$
This upper bound achieves the rate of $\sqrt{T}$, which matches the existing surrogate regret upper bound for bandit multiclass classification in [39]. Regarding the dependence on $K$, our bound improves the existing $O(K\sqrt{T})$ bound in [38, 39] by a factor of $\sqrt{K}$. Note, however, that the $O(K\sqrt{T})$ bound in [39] applies to a broader class of surrogate loss functions. For example, in $K$-class classification, their bound applies to the base-$k$ logistic loss for $k\leq K$, while ours is restricted to the baes-2 logistic loss. A more detailed discussion is given in Appendix C. As for tightness, an $\Omega(\sqrt{T})$ lower bound is provided in [39] for the graph feedback setting, a variant of the bandit feedback model. This suggests that the $\sqrt{T}$ dependence would be close to being tight, although this lower bound does not directly apply to the bandit setting. Therefore, whether the rate of $\sqrt{T}$ is improvable or not is left open.
@@
From the convexity of $S_{t}$ and the unbiasedness of $\widehat{\bm{G}}_{t}$, we have $\mathbb{E}[\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}]\leq\mathbb{E}[\sum_{t=1}^{T}{\langle\widehat{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}].$ By applying Lemma 3.3 and Jensen’s inequality, this term can be further upper bounded by $\mathbb{E}[\sum_{t=1}^{T}{\langle\widehat{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}]\leq\sqrt{2}B\sqrt{\mathbb{E}[\sum_{t=1}^{T}{{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}}]}\leq B\sqrt{\frac{2bK}{q}\mathbb{E}[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}]},$ where we used $\mathbb{E}[\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}]=\frac{\|\bm{G}_{t}\|_{\mathrm{F}}^{2}}{p_{t}(\bm{y}_{t})}\leq\frac{K}{q}\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq\frac{bK}{q}S_{t}(\bm{W}_{t}),$ which follows from $p_{t}(\bm{y})\geq K/q$ and (1). From Assumption 3.2, it holds that $\mathbb{E}[\mathcal{R}_{T}]\leq\mathbb{E}[\sum_{t=1}^{T}{\left((1-a)S_{t}(\bm{W}_{t})-S_{t}(\bm{U})\right)}]+qT\leq B\sqrt{\frac{2bK}{q}\mathbb{E}[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}]}-a\mathbb{E}[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}]+qT\leq\frac{bB^{2}K}{2aq}+qT$, where we used $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$. Plugging $q=B\sqrt{K/T}$ into the last inequality completes the proof.∎
We can also prove a surrogate regret bound of $O(\sqrt{KT\log(1/\delta)}+\log(1/\delta))$, which holds with probabilty $1-\delta$. The precise statement and proof are provided in Section D.2. To prove this high-probability bound, we follow the analysis of Theorem 3.5 and use Bernstein’s inequality. To address the challenges posed by the randomness introduced by bandit feedback, we adopt an approach similar to that used in [39], and arguably, we have simplified their analysis.
3.4 Algorithm based on pseudo-inverse matrix estimator with $O(T^{2/3})$ regret
While the surrogate regret bound of $O(\sqrt{KT})$ achieves the presumably tight dependence on $T$, the dependence on $K=\lvert\mathcal{Y}\rvert$ is undesirable for general structured prediction. In fact, we have $K=\binom{d}{m}$ in multilabel classification with $m$ correct labels and $K=m!$ in ranking with $m$ items. To address this issue, we present an algorithm that significantly improves the dependence on $K$ when the target loss function belongs to a special class of SELF (2) with the following additional assumptions:
Assumption 3.6.
(i) $\bm{V}$ is known and invertible, and $\bm{b}$ and $c(\cdot)$ are known and non-negative. (ii) Let $\bm{Q}=\mathbb{E}_{\bm{y}\sim\mu}[\bm{y}\bm{y}^{\top}]$, where $\mu$ is the uniform distribution over $\mathcal{Y}$. At least one of the following two conditions holds: (ii-a) $\bm{Q}$ is invertible, or (ii-b) for any $\bm{y}\in\mathcal{Y}$, $\bm{V}\bm{y}$ lies in the linear subspace spanned by vectors in $\mathcal{Y}$. (iii) For some $\omega>0$, it holds that $\operatorname*{tr}(\bm{V}^{-1}\bm{Q}^{+}(\bm{V}^{-1})^{\top})\leq\omega.$ (iv) For any $\widehat{\bm{y}}_{t}\in\mathcal{Y}$ and $\bm{y}_{t}\in\mathcal{Y}$, it holds that $|\langle\widehat{\bm{y}}_{t},\bm{V}\bm{y}_{t}\rangle|\leq 1$.
The first and fourth conditions are true in the examples in Section 2.4, assuming that the number of correct labels, $m$, is fixed in multilabel classification. (While the fourth condition does not hold when $m>d/2$ in multilabel classification, we can flip 00 and $1$ in the labels to redefine the problem that satisfies the condition.) The second one holds if $\mathcal{Y}$ consists of $d$ linearly independent vectors or $\bm{V}$ is proportional to the identity matrix; either is true in the examples. Also, deriving reasonable bounds on $\omega$ in those examples is not difficult; see also Section D.4 for details.
Algorithm based on pseudo-inverse matrix estimator
As with Section 3.3, we consider estimating the gradient. Let $\bm{P}_{t}\coloneqq\mathbb{E}_{\bm{y}\sim p_{t}}[\bm{y}\bm{y}^{\top}]$ and define the estimator $\widetilde{\bm{y}}_{t}$ of $\bm{y}_{t}$ by $\widetilde{\bm{y}}_{t}\coloneqq{\bm{V}}^{-1}\bm{P}_{t}^{+}\widehat{\bm{y}}_{t}\langle\widehat{\bm{y}}_{t},\bm{V}\bm{y}_{t}\rangle,$ where $\bm{P}_{t}^{+}$ is the Moore–Penrose pseudo-inverse matrix of $\bm{P}_{t}$. Note that, given that $\bm{b}$ and $c(\cdot)$ are known, we can compute $\langle\widehat{\bm{y}}_{t},\bm{V}\bm{y}_{t}\rangle=L_{t}(\widehat{\bm{y}}_{t})-\langle\widehat{\bm{y}}_{t},\bm{b}\rangle-c(\bm{y}_{t})$. Importantly, $\widetilde{\bm{y}}_{t}$ is unbiased, i.e., $\mathbb{E}_{t}\left[{\widetilde{\bm{y}}_{t}}\right]=\bm{y}_{t}$ from the second requirement of Assumption 3.6.
By using this $\widetilde{\bm{y}}_{t}$, we define the pseudo-inverse matrix estimator $\widetilde{\bm{G}}_{t}$ by $\widetilde{\bm{G}}_{t}\coloneqq\left(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\widetilde{\bm{y}}_{t}\right)\bm{x}_{t}^{\top},$ which is also unbiased, i.e., $\mathbb{E}\big{[}\widetilde{\bm{G}}_{t}\big{]}=\bm{G}_{t}.$ Our estimator is inspired by those used in adversarial linear bandits and adversarial combinatorial full-bandits [14, 1, 8].
We use RDUE with $q=({4\omega B^{2}C_{x}^{2}}/{T})^{1/3}$ as the decoding function (assuming $T\geq 4\omega B^{2}C_{x}^{2}$ for simplicity). To update $\bm{W}_{t}$, we use OGD in Section 3.2 as $\mathsf{ALG}$ with $\eta_{t}={B}/{\sqrt{2\sum_{i=1}^{t}\|\widetilde{\bm{G}}_{i}\|_{\mathrm{F}}^{2}}}.$
Regret bounds
This algorithm achieves the following surrogate regret bound independent of $K$:
@@ 3.4 Algorithm based on pseudo-inverse matrix estimator with $O(T^{2/3})$ regret
The above algorithm achieves $\mathbb{E}[\mathcal{R}_{T}]=O\big{(}\omega^{1/3}T^{2/3}\big{)}.$
The proof can be found in Section D.3. Note that we leverage the structure of OSP when using the pseudo-inverse matrix estimator, which largely differs from the existing approaches to surrogate regret analysis for online classification and OSP [38, 39, 35]. With the pseudo-inverse matrix estimator, we can upper bound the second moment of the gradient estimator $\widetilde{\bm{G}}_{t}$ without $K$, which allows for the surrogate regret bound that does not explicitly involves $K$. This is in contrast to the inverse-weighted gradient estimator in Section 3.3. The inverse-weighted gradient estimator involves division by $p_{t}$, whose lower bound comes from uniform exploration on $\mathcal{Y}$; consequently, its upper bound depends on $K=|\mathcal{Y}|$. In other words, the above pseudo-inverse matrix estimator offers an alternative way to obtain an unbiased gradient estimator while eschewing uniform exploration on $\mathcal{Y}$. However, this comes at the price of a somewhat looser bound on the second moment, which increases the dependence on $T$.
By using Theorem 3.7 as a template, we can derive specific bounds for each problem as follows:
Corollary 3.8.
The above algorithm achieves $\mathbb{E}[\mathcal{R}_{T}]\leq O(d^{2/3}T^{2/3})$ in multiclass classification with the 0-1 loss, $\mathbb{E}[\mathcal{R}_{T}]\leq O\big{(}(d^{5}/m(d-m))^{1/3}T^{2/3}\big{)}$ in multilabel classification with $m$ correct labels and the Hamming loss, and $\mathbb{E}[\mathcal{R}_{T}]\leq O(m^{5/3}T^{2/3})$ in ranking with the number of items $m$ and the Hamming loss.
The proof of Corollary 3.8 is deferred to Section D.4. The bound for multilabel classification with $m$ correct labels can be significantly better than the $O(\sqrt{KT})$ bound in Section 3.3 since $K=\sqrt{\binom{d}{m}}$; similarly, the bound for ranking can be much better than the $O(\sqrt{KT})$ bound since $K=\sqrt{m!}$.
4 Delayed full-information feedback
This section discusses OSP with fixed-delay full-information feedback and provides an algorithm that achieves a surrogate regret bound of $O(\min\{D^{2},D^{2/3}T^{1/3}\})$, a better bound than $O(\sqrt{DT})$ that can be achieved with a standard OCO algorithm under delayed feedback [22]. Below, we make the following assumption based on the randomized decoding of [35].
@@ 4 Delayed full-information feedback
There exists a constant $a\in(0,1)$ that satisfies $\mathbb{E}_{t}\left[{L_{t}(\widehat{\bm{y}}_{t})}\right]\leq(1-a)S_{t}(\bm{W}_{t}).$
@@
We employ the Optimistic Delayed Adaptive FTRL algorithm (ODAFTRL) [17] as $\mathsf{ALG}$, which we detail in Section E.1 for completeness. ODAFTRL computes the linear estimator by $\bm{W}_{t+1}\in\operatorname*{arg\,min}_{\bm{W}\in\mathcal{W}}\left\{\sum_{i=1}^{t-D}\langle\bm{G}_{i},\bm{W}\rangle+\frac{\lambda_{t}\|\bm{W}\|_{\mathrm{F}}^{2}}{2}\right\},$ where $\lambda_{t}\geq 0$ is the regularization parameter. By updating $\lambda_{t}$ using an AdaHedge-type algorithm (AdaHedgeD), ODAFTRL achieves the following regret upper bound:
Lemma 4.2 (Informal version of [17, Theorem 12]).
Consider the delayed full-information feedback setting. Then, for any $\bm{U}\in\mathcal{W}$, ODAFTRL with the AdaHedgeD update of $\lambda_{t}$ achieves a regret bound of $\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}\leq\sum_{t=1}^{T}{\langle\bm{G}_{t},\bm{W}_{t}-\bm{U}\rangle}=O\Big{(}\sqrt{\sum_{t=1}^{T}{(\|\bm{G}_{t}\|_{\mathrm{F}}^{2}+D\|\bm{G}_{t}\|_{\mathrm{F}})}}\Big{)}.$
@@
The above algorithm achieves the following surrogate regret bound:
@@
The above algorithm achieves $\mathbb{E}[\mathcal{R}_{T}]=O(\min\{D^{2},D^{2/3}T^{1/3}\}).$
Notably, this bound is independent of $T$, and we expect that it is fairly close to tight, though no lower bound is known in the prior literature [27]. Note that the proof ideas for deriving surrogate regret bounds in the non-delayed setting [39, 35] differ from those in the standard OCO and multi-armed bandits. This means that we cannot naively extend the analyses of the algorithms for delayed feedback in those settings to our case. Below is the proof sketch, and the complete proof is given in Section E.2.